OpenStudy (anonymous):

find the LCM z^2 - 36 and z^2 + 7z + 6

OpenStudy (anonymous):

Z^2-36=z^2-6^2 =(z-6)(z+6) z^2+7z+6 =(z+1)(z+6) So what's the LCM ?

OpenStudy (anonymous):

Not sure if you're there :) LCM[ (z-6)(z+6) , (z+1)(z+6) ] = (z-6)(z+6)(z+1)

OpenStudy (anonymous):

Let me know if anything is unclear.